OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? @youngstudier @bixslo x + y = -1 x - 3y = 11 A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (anonymous):

@YoungStudier @Bixslo

OpenStudy (anonymous):

@escamer

OpenStudy (anonymous):

@celinegirl

OpenStudy (anonymous):

@west_coast

OpenStudy (escamer):

@Deadly_Roses

OpenStudy (escamer):

she should know this ^-^

OpenStudy (anonymous):

@jere

OpenStudy (anonymous):

@jeremy12345

OpenStudy (anonymous):

@pooja195

OpenStudy (maddy1251):

A.)

OpenStudy (maddy1251):

Simplest way to solve these is just plugging in the answer choices you are given. (x,y) coordinates into your given functions.

OpenStudy (anonymous):

2x - y = -5 x + 3y = 22 A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (anonymous):

@Maddy1251

OpenStudy (anonymous):

Okay let me figure out this question

OpenStudy (maddy1251):

@west_coast you got this?

OpenStudy (anonymous):

Yes ima try a few ways i can do it

OpenStudy (maddy1251):

It's A. Just to let you guys know.

OpenStudy (anonymous):

.-. nevermind

OpenStudy (anonymous):

-3x + 4y = -20 x + 4y = -4 A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (anonymous):

@Maddy1251

OpenStudy (anonymous):

@west_coast

OpenStudy (anonymous):

@gabbyalicorn

OpenStudy (maddy1251):

D.

OpenStudy (anonymous):

Tickets to the movies cost $12 for adults and $8 for children. A group of 14 people went to the movies, and the tickets cost $136. The system of equations models this situation, where x is the number of adults and y is the number of children. x+y = 14 12x + 8y = 136 How many adults and how many children were in the group? A. 9 adults and 5 children B. 6 adults and 8 children C. 4 adults and 10 children D. 8 adults and 6 children

OpenStudy (maddy1251):

One moment.

OpenStudy (maddy1251):

B. The adults are your 'x' functions, the children are your 'y' functions, so just plug in the amount of adults & children into each equation.

OpenStudy (anonymous):

A container holds a combination of nickels and quarters. There are 156 coins in all, and their total value is $21.60. The system of equations models this situation, where x is the number of nickels and y is the number of quarters. x+y = 156 0.05x + 0.25y = 21.6 How many nickels and quarters are in the container? A. 118 nickels and 38 quarters B. 106 nickels and 50 quarters C. 87 nickels and 69 quarters D. 72 nickels and 84 quarters

OpenStudy (anonymous):

@Maddy1251

OpenStudy (maddy1251):

C.

OpenStudy (anonymous):

HOLY JOMOLY YOU'RE SMART! ALL OF THEM WERE CORRECT @Maddy1251

OpenStudy (maddy1251):

@Doll.face what can I say, I love math. :-)

OpenStudy (anonymous):

do you like science?

OpenStudy (anonymous):

@Maddy1251

OpenStudy (maddy1251):

Yup, I excel mostly in Physical, Chemistry, and Biology. I am halfway through my year of Physics.

OpenStudy (anonymous):

reactant → product If there are 4 grams of reactant, how many grams of product are produced by the chemical reaction? A) 0 grams B) 2 grams C) 4 grams D) 8 grams

OpenStudy (anonymous):

@Maddy1251

OpenStudy (maddy1251):

C) 4 Grams. According to the Law of Conservation of Mass, in a chemical reaction, matter is neither created nor destroyed. Because of this the sum of the mass of the reactants will always equal the mass of the products

OpenStudy (anonymous):

A + B + C → D + E Describe the mass of the reactants and products given the balanced chemical reaction. A) The mass of the products is 2/3 that of the reactants. B) The mass of the reactants equals the mass of the products. C) The mass of the reactants is less than the mass of the products. D) The mass of the reactants is greater than the mass of the products.

OpenStudy (anonymous):

@Maddy1251

OpenStudy (maddy1251):

@Doll.face remember what I said yesterday. mass of what you put in (reactants) = mass of what you get out (products)

OpenStudy (anonymous):

c? @Maddy1251

OpenStudy (maddy1251):

No, B. :-) Remember everything should equal reactants = products.

OpenStudy (anonymous):

Matthew gets thirsty and decides to make some fruit punch. He dissolves a packet of the sugar mix into a pitcher of warm water. Which of these best describes what occurred? A) Physical change B) Chemical change C) Both a physical and chemical change D) Neither a physical of chemical change @Maddy1251

OpenStudy (maddy1251):

Remember that a physical change is when no new substances are formed. Chemical changes are. Physical changes can be like, reversed or separated, while a chemical change cannot. Like a rusty nail, you cannot reverse the fact it is rusted. But mixing the packet and water together, you can separate them through evaporation.

OpenStudy (maddy1251):

Next question you ask, can you post in a new thread and mention me in it? This is making my computer lag on OS with all of these posts in one thread. :P

OpenStudy (anonymous):

sure

OpenStudy (maddy1251):

Gracias @Doll.face